Quantitative Aptitude Practice Question and Answer

Q: If 25% of a number is subtracted from a second number, the second number reduces to its four-sixth. What is the ratio of the first number to the second number  ? 2037 0

  • 1
    4:3
    Correct
    Wrong
  • 2
    2:3
    Correct
    Wrong
  • 3
    3:2
    Correct
    Wrong
  • 4
    5:3
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 1. "4:3"
Explanation :

Answer: A) 4:3 Explanation: Let the numbers be p & q q-p/4 = 4q/6 => q-4q/6 = p/4 => 2q/6 = p/4  => p/q = 4/3

Q:

Simplification the following equations shown below

2036 0

  • 1
    1512
    Correct
    Wrong
  • 2
    1521
    Correct
    Wrong
  • 3
    1251
    Correct
    Wrong
  • 4
    1531
    Correct
    Wrong
  • 5
    None of these
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 2. "1521"

Q: 5 + 2 x 10/2 = ? 2034 0

  • 1
    15
    Correct
    Wrong
  • 2
    35
    Correct
    Wrong
  • 3
    10
    Correct
    Wrong
  • 4
    20
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 1. "15"
Explanation :

Answer: A) 15 Explanation: Using BODMAS Rule, First solve division, mutiplication and Addition So here 10/2 = 5 Then 2 x 5 =  10 Now, 5 + 10 = 15.   Hence, 5 + 2 x 10/2 = 15.

Q: A man sell 35 waffers for Rs. 15 each and loses Rs. 40.4. If he sells 35 waffers for Rs. 18, how much does he loss or gain ? 2032 1

  • 1
    Rs. 64
    Correct
    Wrong
  • 2
    Rs. 52
    Correct
    Wrong
  • 3
    Rs. 72
    Correct
    Wrong
  • 4
    Rs. 55
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 1. "Rs. 64"
Explanation :

Answer: A) Rs. 64 Explanation: Given Loss = Rs. 40.4We know loss = cost price - selling price 40.4 = CP - 35(15)=> CP = 40.4 + 525=> CP = Rs. 565.4Now the New selling price of waffer is Rs. 18=> SP = 35(18) = 630Now Gain = SP - CP = 630 -565.4 = Rs. 64.6.

Q: A person walking 6/7 of his usual rate is 30 minutes late. What is his usual time ? 2031 0

  • 1
    2 hrs 50 min
    Correct
    Wrong
  • 2
    3 hrs 25 min
    Correct
    Wrong
  • 3
    3 hrs
    Correct
    Wrong
  • 4
    2 hrs 5 min
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 3. "3 hrs"
Explanation :

Answer: C) 3 hrs Explanation: Let t be the usual time. If the usual rate is 1 unit, then the unusual rate is 6/7.Balance by the distance,1 x t = (6/7)(t + 30)7t = 6(t + 30)we get t as,t = 180 minutes = 3 hrs.

Q: 2500473+ (56)2÷7=? 2030 0

  • 1
    547
    Correct
    Wrong
  • 2
    475
    Correct
    Wrong
  • 3
    455
    Correct
    Wrong
  • 4
    None of these
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 4. "None of these"
Explanation :

Answer: D) None of these Explanation:

Q: In a box, there are four marbles of white color and five marbles of black color. Two marbles are chosen randomly. What is the probability that both are of the same color? 2027 0

  • 1
    2/9
    Correct
    Wrong
  • 2
    5/9
    Correct
    Wrong
  • 3
    4/9
    Correct
    Wrong
  • 4
    0
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 3. "4/9"
Explanation :

Answer: C) 4/9 Explanation: Number of white marbles = 4 Number of Black marbles = 5 Total number of marbles = 9 Number of ways, two marbles picked randomly = 9C2 Now, the required probability of picked marbles are to be of same color = 4C2/9C2  +  5C2/9C2 = 1/6 + 5/18 = 4/9.

Q: A shopkeeper who deals in books sold a book at 16% loss. Had she charged an additional Rs.60 while selling it , her profit would have been 14%. Find the cost price, in rupees, of the book ? 2026 0

  • 1
    Rs. 185
    Correct
    Wrong
  • 2
    Rs. 154
    Correct
    Wrong
  • 3
    Rs. 200
    Correct
    Wrong
  • 4
    Rs. 177
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 3. "Rs. 200"
Explanation :

Answer: C) Rs. 200 Explanation: Let the C.P be 'x'Then, the selling price S.P = x - 16x/100= 84x/100 = 21x/25Now, if the S.P is 60 more, then the profit is 14% => 21x/25 + 60 = x + 14x/100=> 114x/100 - 21x/25 = 60=> (57 - 42)x/50 = 60=> 15x/50 = 60x = 3000/15 = 200 Therefore, the Cost price C.P = x = Rs. 200

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully